# Separating subsets from their images

### v1

@@ 2.3. Reduction to primitive permutation groups
Is the lower bound stated in H asymptotically sharp? If it is not, is $2(\mathbf{m}(K)+\mathbf{m}(H))$ an asymptotically sharp lower bound? Furthermore, by excluding an appropriate infinite family of transitive groups, is it possible to obtain a larger lower bound?
2.4. Primitive permutation groups
Recall that, by Example 5, for $\mathrm{Alt}(n)$ and $\mathrm{Sym}(n)$ in their action on $n$ points, $\mathbf{m}(\mathrm{Alt}(n))=\mathbf{m}(\mathrm{Sym}(n))=\lceil(n+1)/2\rceil$. At the moment, they constitute the the single infinite family of primitive permutation groups where we can prove that a linear asymptotic behavior is exhibited. We suspect that, possibly apart from few well-understood infinite families, the size of the smallest non-self-separable subset is asymptotically proportional with the square root of the degree of the group. This intuition is captured by the following question.
Question 8.
@@ 2.4. Primitive permutation groups
Natural candidates that might exhibit a different behaviour are primitive permutation groups of product action type whose socle is a direct product of alternating groups in their natural action (see M and Lemma 21). If the answer to 8 has a negative answer, then we still propose B, already stated in Section 1, as a possible direction of investigation.
In view of this and the already established square-root lower bound (see C), we focus on obtaining upper bounds for $\mathbf{m}(G)$, and thus dismissing the computation of $\mathbf{f}_{{\mathcal{C}_{{\rm pr}}^{\ast}}}(n)$. Let us start our investigation with an example.
@@
Let $\mathrm{PSL}(d,q)\unlhd G\leq\mathrm{P}\Gamma\mathrm{L}(d,q)$ act on the projective space $\mathrm{PG}(d-1,q)$. To find an upper bound on $\mathbf{m}(G)$, one needs to find a subset $A$ of the projective space such that every image of $A$ under a collineation always intersects $A$. An obvious candidate is a projective subspace of dimension exceeding half of that of the ambient $\mathrm{PG}(d-1,q)$. In this case, we can check that if d1d-1 is even, we can take a subspace of dimension (d1)/2(d-1)/2, and we find that, for large $n$,
@@
This construction is not meaningful for $d=2$, for which we cannot take any projective subspace apart from the line itself. In fact, we know very little of the behaviour of $\mathbf{m}(G)$, where $G\leq\mathrm{Aut}(PG(1,q))$. Furthermore, even for $d\geq 4$, the exponent d/2(d1)d/2(d-1) is always larger than $1/2$ (though it converges to $1/2$ as $d$ goes to infinity), perhaps suggesting that these groups might also provide a negative answer to 8. However, we believe that subsets distinct from subspaces might provide an example of smaller non-self separable set. It would be fascinating if some subgeometry $A$ of $\mathrm{PG}(d-1,q)$ were not self-separable, but all the obvious candidates (such as ovoids or Galois subgeometries) seem to be self-separable or too large.
Similar reasoning can produce a large array of upper bounds for primitive groups of almost simple type in standard action. As in [30], with the term standard actions, we refer to permutation groups that are either alternating, symmetric or classical, and, in the former two cases, the domain consists of $k$-subsets, while in the latter the groups are endowed with a subspace action. (For the list of subspace actions, we refer to [11, Chapter 4].)
@@
$(l)$ $\Omega_{2d}^{\epsilon}(q)$ nonsingular $1$-subspaces, $q$ even $(\heartsuit)$
@@
Table 1. Upper bound for the standard actions of the primitive groups of almost simple type. The asymptotic notation and the symbols $(\spadesuit)$, $(\clubsuit)$ and $(\heartsuit)$ are explained in Remark 10.
Theorem J.
@@
Observe that in the last column of Table 1 the symbols $(\spadesuit)$, $(\clubsuit)$ and $(\heartsuit)$ appear. The first symbol $(\spadesuit)$ appears every time the socle of the classical group into consideration is orthogonal and the action is on nondegenerate parabolic $(2k+1)$-subspaces. In this case, we have a precise value for the asymptotic upper bound computed, which can be read from Table 4. The symbol $(\clubsuit)$ appears whenever we encounter the action of a classical group on totally isotropic or totally singular subspaces. Although we have an explicit upper bound for $\mathbf{m}(G)$, which can be derived from Table 3, the expression are so complicated that we cannot access a precise asymptotic, even if aided by a calculator. Still, some experimentation with these functions shows that, if one fixes $k$ and allows $d$ to grow to infinity, the expected outcome is the asymptotic behavior $\lesssim\sqrt{n}$. Last, the symbol $(\heartsuit)$ appears in case $(l)$: if $k=1$, in all the other cases our method gives trivially an upper bound, and we have not investigated this case by itself to reach a different conclusion.
@@
For most standard actions, for a fixed even value of the parameter $k$ in Table 1, we see that $\mathbf{m}(G)\lesssim\sqrt{n}$. Hence, if the degree is large enough, we can expect $\mathbf{m}(G)$ to be arbitrarily close to the lower bound given in C. Moreover, if $k$ is odd, we can note that, as $k$ gets larger, the upper bound computed in J gets arbitrarily close to $\sqrt{n}$.
Let now move on to the other types of primitive groups as described in the O’Nan–Scott Theorem [29]. In what follows, we adopt the partition into eight classes proposed in [38]. We will recall the properties of the classes we address in Section 8.
@@
The proof of K, provided in Section 8.2, also holds for primitive permutation groups of type holomorph of simple and holomorph of compound, but in these cases the resulting upper bound for $\mathbf{m}(G)$ is meaningless as it exceeds $n$. We leave all the holomorph types out of our analysis.
We now concentrate on primitive groups of type product action and of type compound diagonal. They can be described as being the output of the construction of a primitive wreath product applied to types almost simple and simple diagonal. This fact inspired us to obtain the following general theorem (proved in Section 3.2).
Theorem L.
@@
Therefore, we obtain immediately these corollaries.
Corollary M.
Suppose that there exists a constant $C$ such that 8 has a positive solution for the class $\mathcal{C}_{\rm as}^{\ast}$ of all the primitive permutation groups of type almost simple of non alternating socle. Then, for the infinite class $\mathcal{C}_{\rm pa}^{\ast}$ of all primitive permutation groups of type product action whose socle is not isomorphic to a direct product of alternating groups,
@@
Corollary N.
Suppose that there exists a constant $C$ such that 8 has a positive solution for the class $\mathcal{C}_{\rm sd}$ of all the primitive permutation groups of type simple diagonal with constant $C$. Then, for the infinite class $\mathcal{C}_{\rm cd}$ of all primitive permutation groups of type compound diagonal,
@@
3. Stability of $\mathbf{m}(G)$
In this section, we investigate the stability of $\mathbf{m}(G)$, $\mathbf{f}_{\mathcal{C}}(n)$ and $\mathbf{F}_{\mathcal{C}}(n)$ under certain natural operations applied either to the group $G$ or to the family $\mathcal{C}$.
3.1. Stability with respect to subgroups and subfamilies
Recall that $\mathcal{C}$ is an infinite family of transitive permutation group, and that $\mathcal{C}_{n}$ is the subfamily of permutation groups of degree $n$.
Observe that, for every $G\in\mathcal{C}_{n}$, every subset $A$ of the domain of $G$ satisfying $|A|<\mathbf{f}_{\mathcal{C}}(n)\leq\mathbf{m}(G)$ is self-separable for $G$ by definition. Therefore, to find a lower bound $a(n)+1$ for $\mathbf{f}_{\mathcal{C}}(n)$, we need to prove that, for all $G\in\mathcal{C}_{n}$ and for every subset $A$ in the domain of $G$ such that $|A|\leq a(n)$, $A$ is self-separable for $G$. On the other hand, every subset $A$ of the domain of $G$ with the property that $|A|\geq\mathbf{F}_{\mathcal{C}}(n)\geq\mathbf{m}(G)$ is not self-separable. Hence, to find an upper bound $b(n)+1$ for $\mathbf{F}_{\mathcal{C}}(n)$, we need to prove that, for all $G\in\mathcal{C}_{n}$, there exists a subset $A$ in the domain of $G$ such that $|A|\geq b(n)$ and $A$ is not self-separable for $G$, that is, for every $g\in G$, we can prove that $|A\cap A^{g}|\geq 1$.
@@ 3.1. Stability with respect to subgroups and subfamilies
3.2. Stability with respect to quotients and extensions
@@
Since there is no clear notion of an extension for permutation groups, to mirror Lemma 13, we focus on examining how the parameter of interest relates between a wreath product $H\mathop{\mathrm{wr}}K$ and its constituents $H$ and $K$. We start with the imprimitive action for $H\mathop{\mathrm{wr}}K$. Explicitly, if $\Delta$ is the domain of $H$ and $\Gamma$ of $K$, $H\mathop{\mathrm{wr}}K$ acts on the Cartesian product $\Delta\times\Gamma$ by the law
@@
Proof of H.
Let $G$ be a transitive permutation group that embeds in $H\mathop{\mathrm{wr}}K$ with imprimitive action.
We start by proving the lower bound. Let $A$ be any set of points of size $\mathbf{m}(H)+\mathbf{m}(K)-2$. Either, there exists a block $B$ which contains at least $\mathbf{m}(H)$ points or not. If the latter holds, then, there exists an element $g$ of the base group $\mathrm{Fun}(\Gamma,H)$, such that $A$ and $A^{g}$ intersect trivially. Otherwise, $A$ has a nontrivial intersection with at most $\mathbf{m}(K)-1$ blocks, hence there exists an element $g\in K$ from the top groups that maps the blocks intersecting $A$ to blocks that do not. In any case, we are able to find a permutation $g$ such that $A\cap A^{g}$ is empty, and hence every set of size $\mathbf{m}(H)+\mathbf{m}(K)-2$ is self-separable for $G$. The lower bound immediately follows.
@@
$A=\{(x,y)\mid 1\leq x\leq\mathbf{m}(H),\,1\leq y\leq\mathbf{m}(K)\}\,$,
@@
Proof of L.
Let $R\subseteq\Delta$ be a subset such that, for every $h\in H$, $R\cap R^{h}$ is not empty, and consider
$A=\mathrm{Fun}(\Gamma,R)=\left\{f\in\mathrm{Fun}(\Gamma,\Delta)\mid\forall\gamma\in\Gamma:f(\gamma)\in R\right\}.$
For every $g=(h_{1},h_{2},\ldots,h_{|\Gamma|})k\in H\mathop{\mathrm{wr}}K$, for every $f\in A$ and for every $\gamma\in\Gamma$, we compute
@@
is not empty due to the choice of $R$. In particular, by applying Lemma 12 and by choosing $R$ of cardinality $\mathbf{m}(H)$, we obtain that
@@
3.3. Stability with respect to change of action
We complete the section by establishing a connection between the sought-after parameter for two different actions of the same abstract groups.
@@ 3.3. Stability with respect to change of action
Without loss of generality, we can rearrange the elements of $A$ so that
@@
which is a $k$-subset of $G/K$, for some $k\leq m$. By our choice of $m$, there exists a group element $g\in G$ such that $A^{\prime}\cap A^{\prime}g$ is empty, or, more explicitely, such that
@@
$B^{\prime}=\{Kb_{1},Kb_{2},\dots,Kk_{1}b_{m},Hk_{2}b_{1},\dots,Hk_{|K:H|}b_{m}\}.$
Observe that $|B^{\prime}|\geq k|K:H|$, and that
$\bigcup_{i=1}^{k}Kb_{i}=\bigcup_{i=1}^{k}\left(\bigcup_{j=1}^{|K:H|}Hk_{j}b_{i}\right).$
@@
We would like to point out that Lemma 14 is quite weak in a general setting. For instance, consider the symmetric group $\mathrm{Sym}(n)$ in its natural action (that is, with stabiliser $\mathrm{Sym}(n-1)$) and in its regular action (that is, with a trivial stabiliser). Then, applying the upper bounds from Lemmas 14 and E, we deduce that
@@
The claimed results are obtained simply by taking the minimum and the maximum of the inequalities from Lemma 14 and observing that a linear dilatation of the argument of the function does not change the asymptotic behaviour. ∎

### v2

@@ 2.4. Primitive permutation groups
Natural candidates that might exhibit a different behaviour are primitive permutation groups of product action type whose socle is a direct product of alternating groups in their natural action (see N and Lemma 20). If the answer to 7 has a negative answer, then we still propose B, already stated in Section 1, as a possible direction of investigation.
In view of this and the already established square-root lower bound (see C), we focus on obtaining upper bounds for $\mathbf{m}(G)$, and thus avoiding the computation of $\mathbf{f}_{{\mathcal{C}_{{\rm pr}}^{\ast}}}(n)$. Let us consider the following example.
@@
Let $\mathrm{PSL}_{d}(q)\unlhd G\leq\mathrm{P}\Gamma\mathrm{L}_{d}(q)$ act on the projective space $\mathrm{PG}(d-1,q)$. To find an upper bound on $\mathbf{m}(G)$, one needs to find a subset $A$ of the projective space such that every image of $A$ under a collineation always intersects $A$. An obvious candidate is a projective subspace of dimension at least half of that of the ambient $\mathrm{PG}(d-1,q)$. In this case, we can check that if d1d-1 is even, we can take a subspace of dimension (d1)/2(d-1)/2, and we find that, for large $n$,
@@
This construction is not meaningful for $d=2$, since in that case we cannot take any projective subspace apart from the line itself. In fact, we know very little of the behaviour of $\mathbf{m}(G)$ where $G\leq\mathrm{Aut}(PG(1,q))$ and $q$ is large. Furthermore, even for $d\geq 4$, the exponent d/(2(d1))d/(2(d-1)) is always larger than $1/2$ (though it converges to $1/2$ as $d$ goes to infinity), suggesting that these groups might also provide a negative answer to 7. However, we believe that subsets distinct from subspaces might provide an example of smaller non-self-separable set. It would be fascinating if some subgeometry $A$ of $\mathrm{PG}(d-1,q)$ were not self-separable, but all the obvious candidates (such as ovoids or Galois subgeometries) seem to be self-separable or too large.
Similar reasoning can produce a large array of upper bounds for primitive groups of almost simple type in their standard action. As in [27], the term standard actions, refers to permutation groups that are either alternating, symmetric or classical, and, in the former two cases, the domain consists of $k$-subsets, while in the latter the groups are endowed with a subspace action. (For the list of subspace actions, we refer to [11, Chapter 4].)
@@
$(l)$ $\Omega_{2d}^{\epsilon}(q)$ nonsingular $1$-subspaces, $q$ even $\leq\lfloor\frac{n}{2}\rfloor$
@@
Table 1. Upper bound for the standard actions of the primitive groups of almost simple type. The asymptotic notation and the symbols $(\spadesuit)$ and $(\clubsuit)$ are explained in Remark 9.
Theorem K.
@@
Observe that in the last column of Table 1 the symbols $(\spadesuit)$ and $(\clubsuit)$ appear. The first symbol $(\spadesuit)$ appears every time the socle of the classical group into consideration is orthogonal and the action is on nondegenerate parabolic $(2k+1)$-subspaces. In this case, we have a precise value for the asymptotic upper bound computed, which can be read from Table 3. The symbol $(\clubsuit)$ appears whenever we encounter the action of a classical group on totally isotropic or totally singular subspaces. Although we have an explicit upper bound for $\mathbf{m}(G)$, which can be derived from Table 2, the expression are so complicated that we cannot provide a precise asymptotic, even if aided by a calculator. Still, some experimentation with these functions shows that, if one fixes $k$ and allows $d$ to grow to infinity, the expected outcome is the asymptotic behaviour $\lesssim\sqrt{n}$.
@@
For most standard actions, for a fixed even value of the parameter $k$ in Table 1, we see that $\mathbf{m}(G)\lesssim\sqrt{n}$. Hence, if the degree is large enough, we can expect $\mathbf{m}(G)$ to be arbitrarily close to the lower bound given in C. For these actions, if $k$ is odd, we can note that, as $k$ gets larger, the upper bound computed in K gets arbitrarily close to $\sqrt{n}$. Explicitly, if $s(k,n)$ is the first term of the asymptotic expansion of the upper bounds we have computed, then
$\lim_{k}\frac{s(k,n)}{\sqrt{n}}=1.$
Let us now move on to the other types of primitive groups as described in the O’Nan–Scott Theorem [26]. In what follows, we adopt the partition into eight classes proposed in [35]. We will recall the properties of the classes we address in Section 8.
@@
The proof of L, provided in Section 8.2, also holds for primitive permutation groups of type holomorph of simple and holomorph of compound, but in these cases the resulting upper bound for $\mathbf{m}(G)$ is meaningless as it exceeds $n$. We leave all the holomorph types out of our analysis.
We now concentrate on primitive groups of product action type and of compound diagonal type. They can be described as being the output of the construction of a primitive wreath product applied to almost simple and simple diagonal types, respectively. This fact inspired us to obtain the following general theorem (proved in Section 3.2).
Theorem M.
@@
Therefore, we obtain these corollaries immediately, assuming 7 holds for primitive permutation groups of almost simple type and primitive permutation groups of simple diagonal type, respectively.
Corollary N.
Suppose that there exists a constant $C$ such that the class $\mathcal{C}_{\rm as}^{\ast}$ of all the primitive permutation groups of almost simple type of non alternating socle we have
$\limsup_{n}\frac{\mathbf{F}_{\mathcal{C}_{{\rm pr}}^{\ast}}(n)}{\sqrt{n}}=C.$
Then, for the infinite class $\mathcal{C}_{\rm pa}^{\ast}$ of all primitive permutation groups of type product action whose socle is not isomorphic to a direct product of alternating groups,
@@
Corollary O.
Suppose that there exists a constant $C$ such that the class $\mathcal{C}_{\rm sd}$ of all the primitive permutation groups of type simple diagonal type we have
$\limsup_{n}\frac{\mathbf{F}_{\mathcal{C}_{{\rm pr}}^{\ast}}(n)}{\sqrt{n}}=C.$
Then, for the infinite class $\mathcal{C}_{\rm cd}$ of all primitive permutation groups of type compound diagonal,
@@
3. Properties of $\mathbf{m}(G)$
In this section, we investigate the monotonicity of $\mathbf{m}(G)$, $\mathbf{f}_{\mathcal{C}}(n)$ and $\mathbf{F}_{\mathcal{C}}(n)$ under certain natural operations applied either to the group $G$ or to the family $\mathcal{C}$.
3.1. Monotonicity with respect to subgroups and subfamilies
Recall that $\mathcal{C}$ is an infinite family of transitive permutation groups, and that $\mathcal{C}_{n}$ is the subfamily of permutation groups of degree $n$.
Observe that, for every $G\in\mathcal{C}_{n}$, every subset $A$ of the domain of $G$ satisfying $|A|<\mathbf{f}_{\mathcal{C}}(n)\leq\mathbf{m}(G)$ is self-separable for $G$ by definition. Therefore, to find a lower bound $a(n)+1$ for $\mathbf{f}_{\mathcal{C}}(n)$, we need to prove that, for all $G\in\mathcal{C}_{n}$ and for every subset $A$ in the domain of $G$ such that $|A|\leq a(n)$, $A$ is self-separable for $G$. Similarly, every subset $A$ of the domain of $G$ with the property that $|A|\geq\mathbf{F}_{\mathcal{C}}(n)\geq\mathbf{m}(G)$ is not self-separable. Hence, to find an upper bound $b(n)+1$ for $\mathbf{F}_{\mathcal{C}}(n)$, we need to prove that, for all $G\in\mathcal{C}_{n}$, there exists a subset $A$ in the domain of $G$ such that $|A|\geq b(n)$ and $A$ is not self-separable for $G$, that is, for every $g\in G$, we can prove that $|A\cap A^{g}|\geq 1$.
@@ 3.1. Monotonicity with respect to subgroups and subfamilies
Finally, recall that from Remark 3 that for families $\mathcal{D}\subseteq\mathcal{C}$ and for every positive integer $n\in X_{\mathcal{C}}$,
3.2. Monotonicity with respect to quotients and extensions
@@
In the case of permutation groups, we focus on examining how the parameter of interest relates between a wreath product $H\mathop{\mathrm{wr}}K$ and its constituents $H$ and $K$. We start with the imprimitive action for $H\mathop{\mathrm{wr}}K$. Explicitly, if $\Delta$ is the domain of $H$ and $\Gamma$ of $K$, $H\mathop{\mathrm{wr}}K$ acts on the Cartesian product $\Delta\times\Gamma$ by
@@
Proof of I.
Let $G$ be a transitive permutation group that embeds in $H\mathop{\mathrm{wr}}K$ with imprimitive action, acting on $\Delta\times\Gamma$ as described above. For each element of $\gamma\in\Gamma,$ we will use the notation $\Delta^{\gamma}=\{(\delta,\gamma)|\delta\in\Delta\}.$ Note that the set of disjoint copies of $\Delta$ form a block system for $G,$ and we will refer to these copies as blocks in the proof.
We start by proving the lower bound. Let $A$ be any set of points of size $\mathbf{m}(H)+\mathbf{m}(K)-2$. Either, there exists $\gamma\in\Gamma$ such that $\Delta^{\gamma}\cap A\geq\mathbf{m}(H)$ points or not. If the latter holds, then there is an element $g$ of the base group $\mathrm{Fun}(\Gamma,H)$, such that $A$ and $A^{g}$ intersect trivially. On the other hand, if there is such $\gamma$, then $A$ has a nontrivial intersection with at most $\mathbf{m}(K)-1<\frac{|\Gamma|}{2}$ copies of $\Delta$, hence there exists an element $g\in K$ from the top group that maps the blocks intersecting $A$ to blocks that do not. In any case, we are able to find a permutation $g$ such that $A\cap A^{g}$ is empty, and hence every set of size $\mathbf{m}(H)+\mathbf{m}(K)-2$ is self-separable for $G$. The lower bound immediately follows.
@@
$A=\{(x,y)\mid x\in X\subseteq\Delta,y\in Y\subseteq\Gamma,1\leq X\leq\mathbf{m}(H),\,1\leq Y\leq\mathbf{m}(K)\}\,$,
@@
Proof of M.
Let $R\subseteq\Delta$ be a subset with $|R|=\mathbf{m}(H).$ Hence, for every $h\in H$, $R\cap R^{h}$ is not empty, and consider
$A=\mathrm{Fun}(\Gamma,R)=\left\{f\in\mathrm{Fun}(\Gamma,\Delta)\mid f(\gamma)\in R\right\}\forall\gamma\in\Gamma.$
For every $g=(h_{1},h_{2},\ldots,h_{|\Gamma|})k\in H\mathop{\mathrm{wr}}K$, every $f\in A$ and every $\gamma\in\Gamma$, we see that
@@
is not empty due to the choice of $R$. In particular, by applying Lemma 11 we obtain that
@@
3.3. Monotonicity with respect to change of action
We complete the section by establishing a connection between the self-separability parameter for two different actions of the same abstract groups.
@@ 3.3. Monotonicity with respect to change of action
Without loss of generality, we can define
@@
such that for all $i^{\prime}\in\{1,2,\dots,m\}$, there exists $i\in\{1,2,\dots,k\}$ such that $Ha_{i^{\prime}}\subset Ka_{i}.$ Now $A^{\prime}$ is a $k$-subset of $G/K$, for some $k\leq m$. By our choice of $m$, there exists a group element $g\in G$ such that $A^{\prime}\cap A^{\prime}g$ is empty, or, more explicitly, such that
@@
$B^{\prime}=\{Hk_{1}b_{1},Hk_{1}b_{2},\dots,Hk_{1}b_{m},Hk_{2}b_{1},\dots,Hk_{|K:H|}b_{h}\}.$
Observe that $|B^{\prime}|\geq h|K:H|$, and that
$\bigcup_{i=1}^{h}Kb_{i}=\bigcup_{i=1}^{h}\left(\bigcup_{j=1}^{|K:H|}Hk_{j}b_{i}\right).$
@@
We would like to point out that Lemma 13 is quite weak in a general setting. For instance, consider the symmetric group $\mathrm{Sym}(n)$ in its natural action (that is, with stabiliser $\mathrm{Sym}(n-1)$) and in its regular action (that is, with a trivial stabiliser). Then, applying the upper bounds from Lemmas 13 and F, we deduce that
@@
The claimed results are obtained by taking the minimum and the maximum of the inequalities from Lemma 13 and observing that a linear dilatation of the argument of the function does not change the asymptotic behaviour. ∎